OpenStudy (anonymous):

Eliminate the parameter t. Find a rectangular equation for the plane curve defined by the parametric equations. x = 6 cos t, y = 6 sin t; 0 ≤ t ≤ 2π

OpenStudy (turingtest):

use the fact that\[\sin^2t+\cos^2t=1\]

OpenStudy (anonymous):

i dont even understand how to begin.

OpenStudy (turingtest):

\[x=6\cos t\implies\frac x6=\cos t\]

OpenStudy (anonymous):

so (x/6)^2 + (y/6)^2 = 1 ?

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

so then x^2 + y^2 = 6 ?

OpenStudy (turingtest):

erm, x^2+y^2=6^2

OpenStudy (anonymous):

oooh. thank you

OpenStudy (turingtest):

welcome :)